OpenStudy (anonymous):

ln(x) -ln(4) =6 Please show work

OpenStudy (anonymous):

See: f(x)=lnx^4+e^(4x^2+5) f '(x)=(x^4)'/x^4+e^(4x^2+5).(4x^2+5)' f '(x)=(4x^3)/x^4+e^(4x^2+5).(8x) f '(x)=4/x + 8x.e^(4x^2+5)

OpenStudy (anonymous):

ln(x) = 6 +ln(4) make both sides be raised from e e^(ln(x)) = e^(6+ln(4)) x = e^(6+ln(4)) property of exponents: x = (e^6)(e^ln(4)) x = (e^6)(4) x = 4e^6

OpenStudy (anonymous):

i plugged 4e^6 back into the equation, it works

OpenStudy (anonymous):

\[\ln x=\ln 4+6\]\[x=e^{\ln 4+6}\]\[x=e ^{\ln4}e^6\]\[x=4e^6\]

myininaya (myininaya):

\[\ln(\frac{x}{4})=6\] \[\frac{x}{4}=e^6\] \[x=4e^6\]